OpenStudy (anonymous):

Please help! Perform the operation and leave answers in the form a+bi. (5-2i)/3i a. 2/3-5/3i b. -2/3-5/3i c. -2-5i d. 2-5i

OpenStudy (unklerhaukus):

\[\frac{5-2i}{3i}=\frac{5-2i}{3i}\times1=\frac{5-2i}{3i}\times\frac{-i}{-i}\] \[=\]